Jack and john are both traveling for a family reunion. Jack leaves an hour and a half later. John travels at an average speed of 65 mph. Jack travels at an average speed of 70 mph. How long will it take jack to reach john?
Answer by stanbon(75887)

Jack and john are both traveling for a family reunion. Jack leaves an hour and a half later. John travels at an average speed of 65 mph. Jack travels at an average speed of 70 mph. How long will it take jack to reach john?
---------
Jack DATA:
rate = 70 mph ; time = t-(3/2) hrs ; distance = r*t = 70(t-(3/2)) miles
-----------------
John DATA:
rate = 65 mph ; time = t hrs ; distance = r*t = 65t miles
---------------
Equation:
dist = dist
70t - 105 = 65t
5t = 105
t = 21 hrs
-----------------
t - (3/2) = 19.5 hrs (Jack's time)
-----------------
Cheers,
Stan H.
-----------